how can i find value of sin 135 and all kind of this problems. How can i find values for every angles of every trignometry functions.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
sin x = sin (π – x)
sin 135 = sin (180 – 135) = sin 45 = 1/sqroot(2)
We can also solve this by using the sin (90 + 45)
To find evry angles, we only need to see its combination either in summation form or product form of the angles we know like 30, 45, 60 and 90.
Good Luck!
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty